OpenStudy (anonymous):

Solve |2 - 3x| = 7. Answer {-5/3, 3} {-15, 3} {-3, 5/3} {-3, 15}

jimthompson5910 (jim_thompson5910):

Hint: |2 - 3x| = 7 expands out to 2 - 3x = 7 or 2 - 3x = -7 Let me know if this helps or not. Thanks

OpenStudy (anonymous):

It does thank you!

jimthompson5910 (jim_thompson5910):

you're very welcome

OpenStudy (anonymous):

I lost it